Consider the vectors "\mathbf{a}=(-k, 1+k, k(1+k))" and "\mathbf{b}=(1+k, k(1+k), -k)" where k is some constant. Which of the following is/are correct?



  1. a and b are perpendicular vectors for any k
  2. a∣=∣b∣.
  3. a and b are unit vectors for k=1
  4. The values of k for which ∣a∣=1 are 0 and -1
  5. The two vectors above are parallel.
1
Expert's answer
2021-10-21T14:55:24-0400

1.

If two vectors are perpendicular then "\overrightarrow{A}.\overrightarrow{B}=0"

"[-k\widehat{i}+(i+k)\widehat{j}+9k+k^2)\widehat{k}].[(1+k)\widehat{i}+(k+k^2)\widehat{j}-k\widehat{k}=0"

"-k(1+k)+(1+k)(k+k^2)+(k+k^2)(-k)=0"

"-k-k^2+k+k^2+k^2+k^3-k^2-k^3=0\\0=0"

option 1 is correct


2.

"|\overrightarrow{A}|=\sqrt{(-k)^2+(1+k)^2+(k+k^2)}"


"\\=\sqrt{k^2+1+k^2+2k+k^2+k^4+2k^3}"


"=\sqrt{k^4+2k^3+3k^2+2k+1}"


"\\|\overrightarrow{B}|=\sqrt{(1+k)^2+(k+k^2)+(-k)^2}"


"\\=\sqrt{1+k^2+2k+k^2+k^4+2k^3+k^2}"


"\\\sqrt{k^4+2k^3+3k^2+2k+1}"


"\\|\overrightarrow{A}|=|\overrightarrow{B}|"


option 2 is correct


3.

if "\overrightarrow{A} \& \overrightarrow{B}" is a unit vector


"|\overrightarrow{A}|=1,\space |\overrightarrow{B}|=1"


"k^4+2k^3+3k^2+2k+1=1\\k(k^3+2k^2+3k+2)=0................(1)\\for\space k=1"

equation (1) not satisfied so option (3) is not correct


4.

"|\overrightarrow{A}|=1"

by equation (1)

"k((k^3+2k^2+3k+2)=0\\k=0\\k^3+2k^2+3k+2=0\\put\space k=-1\\-1+2-3+2=0\\0=0"


option 4 is correct


5.

vector is parallel if

"-k=d(1+k)\space \&\space 1+k=d(k(1+k)\\\frac{-k}{1+k}=d,\space \frac{1}{k}=d\space \&\space k(1+k)=d(-k),\space d=(-1+k)"

Therefore the vectors are not parallel